Find the exact values of the six trigonametric functions of theta if theta is in standard position and thepoint P(-3,5) is on the terminal side of theta.
Answer by lwsshak3(11628)

Find the exact values of the six trigonametric functions of theta if theta is in standard position and thepoint P(-3,5) is on the terminal side of theta
**
You are working with a right triangle in Quadrant II with a horizontal leg of -3, vertical leg of 5 and a hypotenuse=√(3^2+5^2)=√(9+25)=√34
Trig functions:
sin theta=5/√34=5√34/34
cos theta=-3/√34=-3√34/34
tan theta=-5/3
csc theta=√34/5
sec theta=-√34/3
cot theta=-3/5